OpenStudy (anonymous):

if f(x)=x^4+4x, find f'(x)

OpenStudy (anonymous):

f'(x)=4x^3+4

OpenStudy (angela210793):

f'(x)=4x^3+4

OpenStudy (angela210793):

(x^n)'=n*x^(n-1) (x)'=1 (c-te)'=0

OpenStudy (anonymous):

can you show me step by step please

OpenStudy (angela210793):

sure (x^4+4x)'=(x^4)'+(4x)'=4*x^(4-3)+4*1=4x^3+4

OpenStudy (anonymous):

ok thanks

OpenStudy (angela210793):

uw :)